# Automatic Forward Model Parameterization with Bayesian Inference of Conformational Populations

### v1

@@ B.1 Methods
Ubiquitin NMR structural ensemble that consists of 10 conformationsCornilescu et al. (1998).
@@
Next, we use an ensemble (144 conformations) from Richter et al. (PDB: 2NR2), where the refinement was performed using the MUMO (minimal under-restraining minimal over-restraining) method Richter et al. (2007). In this approach, simulations started from the X-ray crystal poseVijay-Kumar, Bugg, and Cook (1987). NOEs and $S^{2}$ Lipari–Szabo order parameters were used as NMR structural restraints during a replica-averaged restrained simulation with TIP3P solvent and CHARM22 force field. The augmented potential energy function given as: $E_{total}=E_{CHARMM22}+E_{restraints}$. Scalar couplings were not used during the refinement, but were only used as a validation metric.
@@
Using the Colabfold notebookMirdita et al. (2022) with default parameters, 25 conformations were predicted using RosettaFold2Baek et al. (2023).
@@
An ensemble of conformationa states was derived from a Markov State Model (MSMs) were constructed from a 1-ms simulation of ubiquitin’s native state at 300 K from Piana et al.Piana, Lindorff-Larsen, and Shaw (2013). The PyEMMA Python packageWehmeyer et al. (2019) was used to determine appropriate backbone featurizations using the Variational Approach for Markov Processes (VAMP) scoring function VAMP-2Wu and Noé (2020). Based on the obtained scores, inverse distances were selected as features, while torsions were excluded due to lower VAMP-2 scores and minimal contribution when paired with inverse distances. When comparing distances and inverse distances, the similarity of the average scores and relative standard deviations show that both featurizations are adequate, and inverse distances was selected purely on the higher average scores. Time-lagged independent component analysis (tICA) followed by $k$-means clustering was used to discretize simulations into a structural ensemble of 500 conformational states for MSM construction Pérez-Hernández et al. (2013); J (1967). For the BICePs calculation, each of the 500 states were given equal statistical weight to enforce the uniform prior $p(X)$.
@@
The range of sampled uncertainty parameters $\sigma$ on a grid of logarithmically-spaced values between 0.001 to 100, to enforce the Jeffrey’s prior. Each grid value in the list was a factor of 1.02 larger than the next: [1.00e-03, 1.02e-03, 1.04e-03, 1.06e-03, ... 9.72e+01, 9.92e+01], resulting in a list of 582 values. For the Good-Bad model, sampling of the extra nuisance parameter, $\varphi$ took place on a grid from 1 to 100 with 1000 equally-spaced points.
@@
Optimization was performed for a maximum of 2M steps with a tolerance of 1e-7 and $\alpha=1e-5$. With increasing amounts of data restraint energy, the optimization problem becomes more complicated and more iterations are required to converge. In the case of insufficent iterations, the $\xi$-optimization might return negative $\xi$-values, which is incorrect and not physical. Initially, we start with 11 $\xi$-values {1.0, 0.9, 0.8,...,0.0} and with shift to lower values of $\xi$ after optimization e.g., {1.0, 0.7, 0.56, 0.45, 0.36, 0.28, 0.2, 0.14, 0.08, 0.04, 0.0}.
@@
Figure S2: BICePs predicted forward model parameters in the presence of random and systematic error ($\sigma_{data}=0.471$) for a toy model system. (a) Karplus curves predicted for SVD (orange) and BICePs (blue), where the "true" (black dashed line) parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$. The extracted parameters from the SVD fitting were found to be $\{A=6.11\pm 0.06,B=-1.63\pm 0.04,C=1.80\pm 0.04\}$ and the BICePs was $\{A=6.6\pm 0.037,B=-1.8\pm 0.016,C=1.5\pm 0.027\}$, averaged over three independent chains. The uncertainty is represented by the thickness of the curves. For the BICePs calculation, we used the Good-Bad likelihood model with 32 replicas and burned for 20k steps, followed by 50k steps of additional sampling. (b) The marginal posterior distribution of uncertainty $p(\sigma_{J})$. The maximum a posteriori was determined to be $\hat{\sigma}_{J}=0.272$, and the a posteriori variance scaling parameter $\hat{\varphi}_{J}=1.98$. (c) Landscapes of the BICePs score, $f$ for pairs of Karplus coefficients. (d) The marginal posterior distribution of Karplus coefficients.
Figure S3: The Student’s model gives similar performance to the Good-Bad model. Karplus coefficients predicted using the Student’s likeliood model ($\{A=6.8\pm 0.033,B=-1.9\pm 0.025,C=1.4\pm 0.033\}$) are compared against SVD when faced with random and systematic error ($\sigma_{data}=0.471$). The "True" parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$, the same as Figure S2.
@@
Figure S5: Optimization of $\xi$-values to enhance free energy estimation. The y-axis is thermodyamic length, $|\ell(\xi_{n+1})-\ell(\xi_{n})|$ for each pair of intermediates, derived from the variance in distributions $p(\Delta u_{n,n+1})$. The top panel shows $\xi_{i}$ values before optimization. The bottom panel shows the optimized $\xi_{i}^{*}$ values. Optimized $\xi$-values have shifted to smaller values to make up for the lack of overlap between these adjacent thermodynamic ensembles.
Figure S6: The overlap matrix $O_{ij}$ between thermodynamic states $i$ and $j$ for intermediates along the ($\xi=0\rightarrow 1$) transformation, before optimization of $\xi$-values. The overlap matrix (Klimovich et al 2015) is a Monte Carlo estimate of the average probability of samples in thermodynamic state $j$ being observed in thermodynamic state $i$: $O_{ij}=\left\langle\frac{N_{i}p_{i}(\mathbf{x})}{\sum_{k=1}^{K}N_{k}p_{k}(\mathbf{x})}\right\rangle_{j}$, where $N_{i}$ are the number of samples from thermodynamic state $i$ and $p_{i}(\mathbf{x})$ are posterior probabilities. Significant off-diagonal overlap provides evidence of reliable free energy estimates.
Figure S7: The overlap matrix $O_{ij}$ between thermodynamic states $i$ and $j$ for intermediates along the ($\xi=0\rightarrow 1$) transformation, after optimization of $\xi$-values. Stronger overlap is clearly demonstrated by comparison of post- and pre-optimization (Figure S6).
Figure S8: Variational minimization of the BICePs score can be used to optimize forward model parameters. By variational minimzation of the BICePs score, $f$ optimization traces converge to the "true" parameters. In these calculations, $\xi$-values were optimized prior to running the parameter refinement. We used the Good-Bad model with 4 replicas to minimize computational cost. Karplus coefficients predicted using the Good-Bad likeliood model ($\{6.31\pm 0.02,-1.69\pm 0.03,1.69\pm 0.01\}$) are compared against SVD when faced with random and systematic error ($\sigma_{data}=0.471$). The "True" parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$, the same as Figure S2.

### v2

@@ Methods
This structural ensemble consists of the 10 conformations deposited in the Protein Data Bank (PDB: 1D3Z) from the NMR structural refinement performed by Cornilescu et al. Cornilescu et al. (1998). The ensemble was calculated by the program X-PLOR using 2727 NOE distance restraints and 98 dihedral angle restraints derived from homo- and heteronuclear J couplings.
@@
The 2NR2 structural ensemble (144 conformations) was taken from Richter et al. (PDB: 2NR2), where the refinement was performed using the MUMO (minimal under-restraining minimal over-restraining) method Richter et al. (2007). In this approach, simulations were started from the X-ray crystal poseVijay-Kumar, Bugg, and Cook (1987) in the presence of replica-averaged restraints to NOEs distances and $S^{2}$ Lipari–Szabo order parameters. Simulations used TIP3P solvent and the CHARM22 force field, with an augmented potential energy function $E_{\text{total}}=E_{\text{CHARMM22}}+E_{\text{restraints}}$. Scalar couplings were not used during the refinement, but were only used as a validation metric.
@@
The RF2 structural ensemble was generated using RosettaFold2Baek et al. (2023) made available through a Colabfold notebook.Mirdita et al. (2022) Default parameters were used, and 25 conformations were generated.
@@
The CHARMM22* structural ensemble was derived from a Markov State Model (MSMs) we constructed from a one-millisecond simulation of ubiquitin’s native state at 300 K from Piana et al.Piana, Lindorff-Larsen, and Shaw (2013). The PyEMMA Python packageWehmeyer et al. (2019) was used to determine appropriate backbone featurizations using the Variational Approach for Markov Processes (VAMP) scoring function VAMP-2Wu and Noé (2020). Based on the VAMP-2 scores, inverse distances were selected as features, while torsions were excluded due to lower VAMP-2 scores and their minimal contribution when paired with inverse distances. When comparing distances and inverse distances, the similarity of the average scores and relative standard deviations show that both featurizations are adequate, and inverse distances was selected purely on the higher average scores. Time-lagged independent component analysis (tICA) followed by $k$-means clustering was used to partition the trajectory data into 500 discrete conformational states for MSM construction Pérez-Hernández et al. (2013); J (1967). For the BICePs calculation, each of the 500 states was given equal statistical weight to enforce the uniform prior $p(X)$.
@@
Uncertainty parameters $\sigma$ were sampled on a grid of logarithmically-spaced values between 0.001 to 100, to enforce the Jeffrey’s prior. Each grid value in the list was a factor of 1.02 larger than the next: [1.00e-03, 1.02e-03, 1.04e-03, 1.06e-03, ... 9.72e+01, 9.92e+01], resulting in a list of 582 values. For the Good-Bad model, sampling of the extra nuisance parameter, $\varphi$ took place on a grid from 1 to 100 with 1000 equally-spaced points.
@@
Optimization was performed for a maximum of 2M steps with a tolerance of $10^{-7}$ and $\alpha=10^{-5}$. With increasing amounts of data restraint energy, the optimization problem becomes more complicated and more iterations are required to converge. In the case of insufficent iterations, the $\xi$-optimization might return negative $\xi$-values, which is incorrect and not physical. Initially, we start with 11 $\xi$-values {0.0, 0.1, ... 0.9, 1.0} and with shift to lower values of $\xi$ after optimization e.g., {0.0, 0.02, 0.04, 0.07, 0.12, 0.26, 0.40, 0.60, 0.90, 0.97, 1.0}.
@@
Figure S2: BICePs predicted forward model parameters in the presence of random and systematic error ($\sigma_{data}=0.471$ Hz) for a toy model system. (a) Karplus curves predicted for SVD (orange) and BICePs (blue), where the "true" (black dashed line) parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$. The extracted parameters from the SVD fitting were found to be $\{A=6.11\pm 0.06,B=-1.63\pm 0.04,C=1.80\pm 0.04\}$ and the BICePs was $\{A=6.6\pm 0.037,B=-1.8\pm 0.016,C=1.5\pm 0.027\}$, averaged over three independent chains. The uncertainty is represented by the thickness of the curves. For the BICePs calculation, we used the Good-Bad likelihood model with 32 replicas and burned for 20k steps, followed by 50k steps of additional sampling. Red dots correspond to the syntetic experimental J-coupling data points. (b) The marginal posterior distribution of uncertainty $p(\sigma_{J})$. The maximum a posteriori was determined to be $\hat{\sigma}_{J}=0.272$ Hz, and the a posteriori variance scaling parameter $\hat{\varphi}_{J}=1.98$. (c) Landscapes of the BICePs score, $f$ for pairs of Karplus coefficients. (d) The marginal posterior distribution of Karplus coefficients.
Figure S3: The Student’s model gives similar performance to the Good-Bad model. Karplus coefficients predicted using the Student’s likeliood model ($\{A=6.8\pm 0.033,B=-1.9\pm 0.025,C=1.4\pm 0.033\}$) are compared against SVD when faced with random and systematic error ($\sigma_{data}=0.471$ Hz). The "True" parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$, the same as Figure S2.
@@
Figure S5: Optimization of thermodynamic intermediates $\xi_{k}$ used to calculate the BICePs score, $f_{\xi=0\rightarrow 1}$. In this scenario, sampling of the BICePs energy function $u=-\log p(\mathbf{X},\sigma|D,\theta)$ is performed for a series of intermediates $\xi_{k}$ progressively from $\xi=1$ to $\xi=0$ (a). The samples are used to optimize a new set of $\xi_{k}$ spaced uniformly in thermodynamic length (b). While the thermodynamic overlap matrix for the unoptimized $\xi_{k}$ show poor overlap in the region of low $\xi$ values (c), the overlap matrix for the optimized $\xi_{k}$ shows good overlap for all neighboring intermediates (d). Reproduced from Ref. Novack et al. (2025) with permission.
Figure S6: Variational minimization of the BICePs score can be used to optimize forward model parameters. By variational minimzation of the BICePs score, $f$ optimization traces converge to the "true" parameters. In these calculations, $\xi$-values were optimized prior to running the parameter refinement. We used the Good-Bad model with 4 replicas to minimize computational cost. Karplus coefficients predicted using the Good-Bad likeliood model ($\{6.31\pm 0.02,-1.69\pm 0.03,1.69\pm 0.01\}$) are compared against SVD when faced with random and systematic error ($\sigma_{data}=0.471$ Hz). The "True" parameters were set to be $\{A=6.51,B=-1.76,C=1.6\}$, the same as Figure S2.
@@
Figure S16: BICePs trains NN parameters for $J$-coupling predictions using a toy model system. (a) Minimization of the BICePs score during training of five independent neural networks (NNs). The solid blue line indicates the average BICePs score across the chains, with the shaded region representing the standard deviation. Convergence is evident by the plateau in score near 1500 training steps. (b) Karplus curves predicted by each of the five independently trained NNs. The mean prediction across all networks is shown in blue with uncertainty indicated by curve thickness (standard deviation across networks). The true curve (black dashed) corresponds to the parameters ${A=6.51,B=-1.76,C=1.60}$. Predictions obtained by singular value decomposition (SVD) are shown in orange, with extracted parameters ${A=6.47\pm 0.002,B=-1.75\pm 0.001,C=1.62\pm 0.001}$. BICePs calculations were performed using a Good-Bad likelihood model and 32 replicas. No random or systematic noise was added to the synthetic data. Red dots correspond to the syntetic experimental J-coupling data points.
Figure S17: BICePs trains NN parameters for $J$-coupling predictions using a toy model system in the presence of random and systematic error ($\sigma_{\text{data}}=0.79$ Hz). (a) Minimization of the BICePs score during training of five independent neural networks (NNs). The solid blue line indicates the average BICePs score across the chains, with the shaded region representing the standard deviation. Convergence is evident by the plateau in score near 1000 training steps. (b) The average gradient norm across the five chains during training. (c) Karplus curves predicted by each of the five independently trained NNs. The mean prediction across all networks is shown in blue with uncertainty indicated by curve thickness (standard deviation across networks). The true curve (black dashed) corresponds to the parameters ${A=6.51,B=-1.76,C=1.60}$. Predictions obtained by singular value decomposition (SVD) are shown in orange, with extracted parameters ${A=6.39\pm 0.13,B=-1.55\pm 0.06,C=1.94\pm 0.10}$. BICePs calculations were performed using a Good-Bad likelihood model and 32 replicas. Red dots correspond to the syntetic experimental $J$-coupling data points.